# Learning Logical Rules using Minimum Message Length

### v1

@@ Introduction
Equation 2 is the MML length function, which balances hypothesis complexity and data fit. A complex hypothesis must explain the data succinctly to be selected over a simple hypothesis that less succinctly explains the data. An MML optimal hypothesis is one that has the smallest value of $C(H,E)$.
We now describe how to define the message lengths for each part of Equation 2.
Hypothesis Message Length: $C(H)$
We now describe the type of hypothesis we communicate.
Probabilistic Logic Programs
We learn a set of rules $R$ that generalise the relation p/a from training examples. We use these rules to predict the probability that an example is true or false. We describe each case with a separate probabilistic rule with Problog semantics (De Raedt, Kimmig, and Toivonen 2007):
Case 1: When $R$ entails an example, the example is true with probability $\theta_{+}$ and false with probability $1-\theta_{+}$, described with the probabilistic rule:
$\theta_{+}::\texttt{predTrue(}X_{1},\dots,X_{a}\texttt{):-p(}X_{1},...,X_{a}\texttt{).}$
An example is predicted to be true when the probability of $\texttt{predTrue(}X_{1},\dots,X_{a}\texttt{)}$ is greater than $0.5$.
Case 2: When $R$ does not entail an example, the example is false with probability $\theta_{-}$ and true with probability $1-\theta_{-}$, described with the probabilistic rule:
$\theta_{-}::\texttt{predFalse(}X_{1},\dots,X_{a}\texttt{):-not }\texttt{p(}X_{1},\dots,X_{a}\texttt{).}$
We need two probabilistic rules because cases 1 and 2 are mutually exclusive, so they must be described separately. As these cases also cover all example possibilities, the union of their respective rules provides a probabilistic prediction for any example. We take the union of these rules with $R$ as our MML hypothesis $H$. We denote the probabilistic rules corresponding to cases 1 and 2 of the hypothesis $H$ as $H_{+}$ and $H_{-}$, respectively. We only send $R$, $\theta_{+}$, and $\theta_{-}$ in the message and assume the receiver knows how to construct $H$ from this information.
Example Group Size Length:
Each sub-hypothesis $H_{+}$ and $H_{-}$ communicates the examples that they respectively entail. Different sub-hypotheses entail different training examples, so partition the example set into two disjoint subsets. $H_{+}$ communicates the true positives and false positives. $H_{-}$ communicates the true negatives and false negatives. These subsets are sent separately, so their sizes must be communicated to ensure that the receiver knows how many examples to expect per set. We define $tp(H)$, $fp(H)$, $tn(H)$, $fn(H)$ as the total number of true positives, false positives, true negatives and false negatives predicted by the hypothesis $H$ on the training examples, respectively. Let $N$ be the total number of training examples known a priori to the receiver. The sender sends the total number of positive predictions $k=tp(H)+fp(H)$. The receiver infers the number of negative predictions $tn(H)+fn(H)$ as NkN-k. The length of communicating a particular $k$ value is based on its probability, which we model as the outcome of two independent binomial trials. These trials have a known prior probability of classification error $r$ and result in hypothetical classifications of the training examples into $tp$, fpfp, tntn, fnfn true positives, false positives, true negatives and false negatives, respectively.
  1. 1.
    Trial 1: classifying the $tp+fn$, positive examples. A trial of $tp+fn$ flips, where each success corresponds to a $tp$ and each failure, an $fn$.
  2. 2.
    Trial 2: classifying the $tn+fp$, negative examples. A trial of $tn+fp$ flips, where each success corresponds to a $tn$ and each failure, an fpfp
Then the probability of a specific, hypothetical $tp,fp$ outcome is the product of the probabilities of these two trials:
$P(tp,fp)$ $=\binom{tp+fn}{tp}(1-r)^{tp}r^{fn}$ (3)
$\times\binom{tn+fp}{tn}(1-r)^{tn}r^{fp}$
The probability of $k$ positive predictions is the sum of the probabilities of all $(tp,fp)$ outcomes such that $tp+fp=k$. The length of sending $k$ is the negative log of this probability.
Hypothesis Syntax Length: $C_{R}(H)$
The receiver can deduce the complete hypothesis $H$ from (i) the set of rules $R$ that generalise the examples and (ii) the probabilities with which each subhypothesis holds ($\theta_{+}$ and $\theta_{-}$). We define three functions which map a hypothesis $H$ to these components: $\mathbf{R}(H)=R$, $\mathbf{\Theta_{+}}(H)=\theta_{+}$ and $\mathbf{\Theta_{-}}(H)=\theta_{-}$. We first describe the communication of the syntax of $R$, which we encode piecewise. We use the example of communicating the syntax of the rule:
         {p(X,Y):-p1(X,Y), p2(X).}
We assume that all variables in the head literal are distinct and that the head literal is already known to the receiver from the task. Therefore, only body literals need to be communicated. Communication of $R$ begins by stating the total number of body literals $l(R)$ in the set of rules, where $l(R)\in\{0,1,\dots,M\}$ and $M$ is the maximum allowed number of literals. We assume no literal number is more likely than any other, so we use a uniform code over the possible $l(R)$ values and thus ignore it in the message length. After specifying the total number of body literals $l(R)$, the sender communicates how they are grouped into rules. The number of possible literal groupings is given by the integer partition function $partition(l(R))$. In our example, there are two possible groupings:
    1) p(X,Y):-?,?.    2) p(X,Y):-?.
                          p(X,Y):-?.
This part of the message has length $\log{partition(l)}$ and is larger for programs with more literals.
After sending the literal groupings, the sender sends each rule $r\in R$ separately. The sender first communicates the predicate symbols of the literals in the body of $r$. To communicate the predicate symbols, we model their probability. The receiver knows a priori the probability of seeing a predicate symbol. We calculate these probabilities using the Herbrand Base, which is the set of all literals that can be constructed from the examples and background knowledge in a given task that is consistent with the given syntactic bias (Nienhuys-Cheng and de Wolf 1997). Let $o_{p}$ be the number of atoms that contain the predicate symbol $p$ in a Herbrand Base $HB$. Then the probability of the predicate symbol $p$ is $P_{pred}(p)=\frac{o_{p}}{|HB|}$. Therefore, more general predicate symbols have greater probabilities. We only consider non-recursive programs, so we omit the counts of head predicates in the computation of $P_{pred}(p)$. Let $ps(r)$ be the predicate symbols in the body of the rule $r$ and $fr(p,r)$ be the number of times the predicate symbol $p$ appears in the body of $r$. Then the probability of seeing the predicate symbols in $r$ is the multinomial formula:
$P_{preds}(r)=\frac{\left[\sum_{p\in ps(r)}\text{fr}(p,r)\right]!}{\prod_{p\in ps(r)}\text{fr}(p,r)!}\prod_{p\in ps(r)}\!\!\!P_{\text{pred}}(p)^{\text{fr}(p,r)}$ (4)
The message length required to communicate the predicate symbols in a rule is $\log{P_{preds}(r)}$.
After sending the predicate symbols, the sender sends the body variables. The sender first sends the frequency of each distinct variable in the rule. Let $\text{fr}(v_{i},r)$ be the number of times the variable $v_{i}$ appears in rule $r$. We define var_slots_body(r)var\_slots\_body(r) as the total arity of all the body literals in rule $r$ and d_vars(r)d\_vars(r) as the number of distinct variables in $r$. Then the number of valid $\text{fr}(v_{i},r)$ assignments to each of the d_vars(r)d\_vars(r) distinct variables in the rule, n_freqs(r)n\_freqs(r), is given by stars and bars (Ross 1976):
$n\_freqs(r)=\left(\begin{smallmatrix}d\_vars(r)+var\_slots\_body(r)-1\\ d\_vars(r)-1\end{smallmatrix}\right)$ (5)
The sender specifies the variable frequencies with a message of length $\log{n\_freqs(r)}$.
To complete the rule, the sender communicates the specific sequence of variables that fill the empty body variable slots. The permutations of identical variables or relabelling of body-only variables (such as Y and Z below) produce semantically equivalent programs.
 h(X):-p1(Y),p2(Z). = h(X):-p1(Z),p2(Y).
Accounting for equivalence, the number of possible variable sequences n_seqs(r)n\_seqs(r) in a rule with d_body_vars(r)d\_body\_vars(r) distinct body-only variables, where variable $j$ is repeated $\text{fr}(v_{j},r)$ times is:
$n\_seqs(r)=\frac{var\_slots\_body(r)!}{d\_body\_vars(r)!\prod_{i}^{d\_vars(r)}\text{fr}(v_{i},r)!}$
The sender specifies a single variable sequence with a message of length $\log{n\_seqs(r)}$. Then the message length required to communicate $R$ is:
$C_{R}(H)=\log{partition(l(\mathbf{R}(H)))}+C_{R^{\prime}}(H)$ (6)
$C_{R^{\prime}}(H)=\sum_{r\in\mathbf{R}(H)}\Big{[}$- $\log{P_{preds}(r)}+\log{n\_seqs(r)}$ (7)
$+\log{n\_freqs(r)}\Big{]}$
It is convenient to delineate $C_{R^{\prime}}(H)$ as it has additive linearity properties, which we later use in our optimisation.
Theta Length: $C_{\theta_{+},\theta_{-}}(H)$
We describe communication of the $\theta_{+}=\mathbf{\Theta}_{+}(H)$ and $\theta_{-}=\mathbf{\Theta}_{-}(H)$ values, the Bernoulli probabilities with which $H_{+}$ and $H_{-}$ each hold. An MML approximation (Wallace and Freeman 1987) to the message length of communicating these values is known. We state this length and refer the reader to the full derivation in Wallace (2005). Let $\alpha$ and $\beta$ represent the parameters of the beta prior distribution $\mathcal{B}(\alpha,\beta)$ over each $\theta$ parameter. Then the length of communicating $\theta_{+}$ is $C_{\theta_{+}}(H)$:
$(\alpha-\frac{1}{2})\log{\mathbf{\Theta}_{+}(H)}-(\beta-\frac{1}{2})\log{\big{[}1-\mathbf{\Theta}_{+}(H)\big{]}}+K$ (8)
Where $K\approx-\log{\mathcal{B}(\alpha,\beta)}-0.7425+\log{\binom{tp(H)+fp(H)}{tp(H)}}$, and:
$\mathbf{\Theta}_{+}(H)=\frac{tp(H)+\alpha-0.5}{tp(H)+fp(H)+\alpha+\beta-1}$ (9)
Similarly, the length of communicating $\theta_{-}$ is $C_{\theta_{-}}(H)$:
$(\alpha-\frac{1}{2})\log{\mathbf{\Theta}_{-}(H)}-(\beta-\frac{1}{2})\log{\big{[}1-\mathbf{\Theta}_{-}(H)\big{]}}+K$ (10)
Where $K\approx-\log{\mathcal{B}(\alpha,\beta)}-0.7425+\log{\binom{tn(H)+fn(H)}{tn(H)}}$, and:
$\mathbf{\Theta}_{-}(H)=\frac{tn(H)+\alpha-0.5}{tn(H)+fn(H)+\alpha+\beta-1}$ (11)
The message length required to communicate the theta values is the sum of the message lengths required to communicate $\theta_{+}$ and $\theta_{-}$:
$C_{\theta_{\pm}}(H)=C_{\theta_{+}}(H)+C_{\theta_{-}}(H)$ (12)
Total Hypothesis Length
The total message length of the hypothesis $C(H)$ is the sum of equations 6 and 12, and the length of communicating the sizes of the example sets:
@@
Where $k$ is the size of the group communicated by $H_{+}$.
Examples Length: $C(E|H)$
We describe how to send examples $E$ using the hypothesis $H$. We define a likelihood function $P(E|H)$ which determines $C(E|H)$ (Equation 1). Each example is composed of a literal and a truth value. We describe sending each in turn.
Literals
We assume that no particular example is a priori more likely to be seen than any other. Let $e(H_{i})$ be all literals in the Herbrand Base entailed by the subhypothesis $H_{i}$, $i\in\{+,-\}$. Then the probability of observing a particular set of examples is the chance of randomly selecting $tp(H)+fp(H)$ examples from $e(H_{+})$ and $tn(H)+fn(H)$ examples from $e(H_{-})$:
$P_{lits}(E|H)=\left[\binom{|e(H_{+})|}{tp(H)\!+\!fp(H)}\binom{|e(H_{-})|}{tn(H)\!+\!fn(H)}\right]^{-1}$
The length of this message is $\log{P_{lits}(E|H)}$.
Truth Values
According to Problog semantics, the probability of a sequence of $tp(H)$, $fp(H)$, $tn(H)$, and $fn(H)$ examples is given by the product of two Bernoulli distributions:
$P_{T/F}(E|H)=\theta_{+}^{tp(H)}(1-\theta_{+})^{fp(H)}\times\theta_{-}^{tn(H)}(1-\theta_{-})^{fn(H)}$
The length of this message is $\log{P_{T/F}(E|H)}$
Total
The length of the message required to communicate the examples is therefore:
$C(E|H)=-\log{P_{lits}(E|H)}-\log{P_{T/F}(E|H)}$ (14)
Total Message Length
The receiver now has all the necessary information to reconstruct both the communicated hypothesis and data. The full message length is the sum of Equations 13 and 14. We search for the hypothesis $H$ which minimises $C(H,E)$, our cost function.
Algorithms
@@
To test claims 1-6, we need algorithms that (i) compare exact MML and MDL programs and (ii) efficiently learn programs with good MML and MDL scores. Our contribution in this section is two such algorithms: RandomLearner and ApproxLearner.
RandomLearner
@@
As the exact MML score is difficult to optimise, this learner calculates exact MML and MDL scores for a large number of random programs, selecting programs with the smallest MML and MDL scores. We generate $T$ random programs for each task. The method of generating the random programs follows. For each size 1Max_Size1-Max\_Size, we generate a maximum of $I$ rules of this size which conform to the syntactic bias given the task. Rules are constrained to only contain at most v_maxv\_max variables. From these rules, we generate a random program. We choose a random number $c$ between 1 and Max_ClausesMax\_Clauses, and randomly select $c$ rules from the set of all generated rules. This random set of rules is a random program. We repeat this $T$ times to generate all random programs. Specific values for T,Max_Size,I,v_max,c,T,\ Max\_Size,\ I,\ v\_max,\ c, and Max_ClausesMax\_Clauses may be found in Section 6.
ApproxLearner

### v2

@@ 4 Minimum Message Length
Equation 1 is the MML cost function, which balances hypothesis complexity and data fit. A complex hypothesis must explain the data succinctly to be selected over a simple hypothesis that less succinctly explains the data. An MML optimal hypothesis is one that has the smallest value of $C(h,E)$.
Equation 1 is formed of a hypothesis cost $C(h)$ and an examples cost $C(E|h)$. We describe each at a high-level. Full derivations for omitted equations may be found in the Appendix.
4.1 Hypothesis Cost
The cost of sending a hypothesis $h$ (definition 2) is:
$C(h)=\text{syntax}(h)+\text{theta}(h)+\text{coverage}(h)$ (2)
We discuss each component in turn.
4.1.1 Syntax
The cost of the syntax of $h$ is the sum of three subcosts [^1] We omit sending the total number of literals in $h$, as its cost is constant for any hypothesis:
$\text{syntax}(h)=\text{structure}(h)+\sum_{r\in\text{det}(h)}\big[\text{predicates}(r)+\text{vars}(r)\big]$
We discuss each subcost in turn.
Structure.
The cost of the arrangement of literals into rules in $\text{det}(h)$, denoted $\text{structure}(h)$, is based on the total number of ways that the non-head literals can be arranged into rules, which is the integer partition function. The sender specifies a single arrangement. For instance, if there are $3$ total non-head literals in $\text{det}(h)$, there are $3$ possible arrangements:
 f(X):-*,*,*.  f(X):-*,*.  f(X):-*.
               f(X):-*.    f(X):-*.
                           f(X):-*.
Specifying one arrangement from the above three requires $\log{3}$ bits. Formally, let $\text{body}(r)$ return the literals in the body of rule $r$ and $\text{part}(n)$ denote the integer partition of $n$. Then:
$\text{structure}(h)=\log{\text{part}\left(\sum_{r\in\text{det}(h)}|\text{body}(r)|\right)}$ (3)
The cost is greater for hypotheses with more non-head literals.
Predicate Symbols.
The cost of the predicate symbols in a rule $r$, denoted $\text{predicates}(r)$, is based on the individual predicate generalities, defined as the relative occurrence of the predicate in the Herbrand Base defined over $\text{det}(h)\cup B$ (see definition 1). Formally, let $o_{p}$ be the number of atoms that contain the predicate symbol $p$ in a Herbrand Base $HB$ and $ord(r)$ denote the number of possible orderings of the predicates in the rule. For example, given $r=f(X,Y)\leftarrow a(X),b(Y).$ there are $ord(r)=2$ orderings: $[a/1,b/1]$ and $[b/1,a/1]$. Let $\text{sym}(r)$ return the multiset of the predicate symbols in $r$. Then
$\text{predicates}(r)=-\log{ord(r)}-\sum_{p\in\text{sym}(r)}{\log\frac{o_{p}}{|HB|}}$ (4)
More general (larger $o_{p}$) predicate symbols result in a shorter message length. We only consider non-recursive programs, so we omit the counts of head predicates in the computation of $HB$.
Variables.
The cost of the variables in a rule, denoted $\text{vars}(r)$, is based on the number of possible variable arrangements in the rule. Formally, let $\text{var\_comb}(r)$ represent the number of possible variable combinations in rule $r$, given a set of predicate symbols and a maximum variable count. Then
$\text{vars}(r)={\log{\text{var\_comb}(r)}}$ (5)
4.1.2 Theta Values
The cost of the theta values $(\theta_{+},\theta_{-})$ is based on a standard MML code (MML87) for binomial values Wallace (2005); Wallace and Freeman (1987), denoted as $\text{theta}(h)$ (see Appendix). The MML87 code is shorter for theta values that are similar to our prior beliefs on likely values of $(\theta_{+},\theta_{-})$.
4.1.3 Coverage
We also send the number of entailed examples (definition 6) denoted $N_{+}$, and the number of non-entailed examples, denoted $N_{-}$. These numbers are necessary to decode the message as the two cases are governed by two separate probability models (rules $r_{+}$ and $r_{-}$, see definition 2). The receiver must know how many examples to decode using each probability model. If $h$ makes no errors, then $N_{+}=|E^{+}|$ and $N_{-}=|E^{-}|$. The more errors made by $h$, the larger the difference between $N_{+}$ and $|E^{+}|$. We expect accurate $h$, so state larger differences with longer messages (see Appendix).
4.2 Examples cost
The cost of sending the examples $E$ is:
$C(E|h)=\text{atoms}(E|h)+\text{labels}(E|h)$ (6)
The receiver knows the set of possible examples (the instance space), but not the specific training examples or their truth labels which are sent separately. We describe each next.
Atoms.
We use the terms entailed atoms, non-entailed atoms, entailed examples and non-entailed examples as defined in definition 6. We assume $N_{+}=tp(h)+fp(h)$ example atoms are chosen from the entailed atoms, and $N_{-}=tn(h)+fn(h)$ from the non-entailed atoms. Then the cost of sending the example atoms, denoted $\text{atoms}(E|h)$, is based on two things. First, the probability of choosing the entailed examples. Second, the probability of choosing the non-entailed examples. Then $\text{atoms}(E|h)$ is the sum of the negative logarithm of the two probabilities. Formally, let $e(h)$ and $\neg e(h)$ denote the number of entailed atoms and non-entailed atoms respectively. Let $h$ have $tp(h)$ true positives, $fp(h)$ false positives, $tn(h)$ true negatives, and $fn(h)$ false negatives on $E$. Assuming all $E$ are equally likely to be chosen, $\text{atoms}(E|h)$ is:
$\underbrace{-\log{\binom{|e(h)|}{tp(h)\!+\!fp(h)}^{-1}}}_{\text{ entailed by det}(h)}-\underbrace{\log{\binom{|\neg e(h)|}{tn(h)\!+\!fn(h)}^{-1}}}_{\text{not entailed by det}(h)}$
More probable choices result in a shorter message.
Labels.
The cost of the example truth labels, denoted $\text{labels}(E|h)$, is determined by their probability given $h$ under Problog semantics. The probability of each example truth label is determined by either $r_{+}$ or $r_{-}$ (definition 2). The probability of all example truth labels is the product of the individual truth labels: the product of two Bernouilli distributions which describe the entailed and non-entailed examples respectively (definition 6). Then $\text{labels}(E|h)$ is the sum of the negative logarithms of the two probabilities.
$\underbrace{-\log{\left[\theta_{+}^{tp(h)}(1-\theta_{+})^{fp(h)}\right]}}_{\text{ entailed by det}(h)}-\underbrace{\log{\left[\theta_{-}^{tn(h)}(1-\theta_{-})^{fn(h)}\right]}}_{\text{not entailed by det}(h)}$
The $\text{labels}(E|h)$ encoding favours hypotheses that misclassify fewer examples.
5 MML Example
Consider the task of learning an integer concept $f/1$. Assume we have the background knowledge:
@@ 5 MML Example
Examples $E=E^{+}\cup E^{-}$, where:
$E^{+}=\{f(2),f(5)\}$ $E^{-}=\{f(3)\}$
Consider the hypothesis $h=\{g_{\phi},r_{+},r_{-}\}$ where:
$g_{\phi}=\phi(X)\leftarrow prime(X).$
$r_{+}=0.88::f(X)\leftarrow\phi(X).$
$r_{-}=0.14::f(X)\leftarrow\text{not }\phi(X).$
From definition 2, the deterministic counterpart of $h$ is:
$\text{det}(h)=f(X)\leftarrow prime(X).$
5.0.1 Determining the Theta Values
We calculate the $\theta_{+},\theta_{-}$ values of $h$ from $\text{det}(h)$, applying the MML87 procedure to the results of $\text{det}(h)$ on training examples: $tp(h)=2$, $fp(h)=1$, $tn(h)=0$, $fn(h)=0$. Assuming a beta prior with parameters $\alpha=10$, $\beta=1$ for both $\theta_{+}$ and $\theta_{-}$, MML87 gives $\theta_{+}=0.88$ and $\theta_{-}=0.86$. Note $1-\theta_{-}=0.14$.
Hypothesis Cost
We calculate the costs of $C(h)$ (equation 2).
Structure.
Given that there is only one non-head literal in $h$, there is only one arrangement of literals into rules and thus $\text{structure}(h)=\log{\text{part}(1)}=\log{1}=0$ bits (equation 3).
Predicate symbols.
There is only one rule $r\in\text{det}(h)$ with a single body predicate, $prime/1$. The Herbrand Base, excluding the possible examples, is just the background knowledge. There are $o_{p}=4$ literals which contain the $prime/1$ symbol out of a total $|HB|=16$ (44 $prime/1$, 66 $even/1$, 66 $odd/1$ literals). The cost of $prime/1$ is thus $\text{predicates(r)}-\log{\frac{4}{16}}=2$ bits (equation 4).
Variables.
One single-argument predicate has only one variable arrangement, so $\text{vars}(r)=\log{1}=0$ bits (equation 5).
Theta values.
Given the beta prior, the cost of sending the theta values $\theta_{+}=0.88,\theta_{-}=0.86$ using MML87 is $\text{theta}(h)=6.5$ bits.
Coverage.
The program $\text{det}(h)$ entails all three observed examples in $E^{+}$ and $E^{-}$, therefore, $N_{+}=3$. Assuming a prior misclassification probability of $0.1$, the probability that $P(N_{+}=3)=0.081$. The cost of stating $N_{+}$ is $\text{coverage}(h)=-\log{0.081}\approx 3.6$ bits.
Total hypothesis cost.
The total syntax cost $\text{syntax}(h)$ is the sum of $\text{structure}(h)$, $\text{predicates}(r)$ and $\text{vars}(h)$: $\text{syntax}(h)=0+2+0=2$ bits. The total hypothesis cost (equation 2) is therefore $C(h)=2+6.5+3.6=12.1$ bits.
Examples Cost
We now describe communicating the examples under $h$.
Atoms.
The deterministic hypothesis $\text{det}(h)$ entails all three training examples, and $4$ out of the $10$ possible examples $\{f(0),f(1),f(2),...f(9)\}$. The probability of choosing the $3$ literals in $E$ from the $4$ possible examples is $\binom{4}{3}^{-1}=0.25$, so $\text{atoms}(E|h)=-\log{0.25}=2$ bits.
Labels.
The hypothesis $h$ classifies two examples correctly and one incorrectly under the $\theta_{+}$ probability model. The probability of making these errors is: $0.88^{2}\times(1-0.88)=0.093$ so thus $\text{labels}(E|h)=-\log{0.093}=3.4$ bits.
Total cost.
The total example length is therefore $2+3.4=5.4$ bits.
5.1 Total Length
The total length (equation 1) is $12.1+5.4=17.5$ bits. We search for the $\text{det}(h)$ which minimises the total length.
6 Algorithms
@@
To test claims 1-5, we need algorithms that (i) compare exact MML and MDL hypotheses and (ii) efficiently learn hypotheses with good MML and MDL scores. Our contribution in this section is two such algorithms: RandomLearner and ApproxLearner.
6.1 RandomLearner
@@
As the exact MML score is difficult to optimise, this learner calculates exact MML and MDL scores for a large number of random hypotheses, selecting hypotheses with the smallest MML and MDL scores. We generate $T$ random hypotheses for each task. The method of generating the random hypotheses follows. For each size 1Max_Size1-Max\_Size, we generate a maximum of $I$ deterministic rules of this size which conform to the syntactic bias given the task. Rules are constrained to only contain at most v_maxv\_max variables. From these rules, we generate a random, deterministic program, $\text{det}(h)$. We choose a random number $c$ between 1 and Max_ClausesMax\_Clauses, and randomly select $c$ rules from the set of all generated rules. This random set of rules is a random program. We repeat this $T$ times to generate all random programs. We convert all programs $\text{det}(h)$ to probabilistic hypotheses $h$, scoring each as described. Specific values for T,Max_Size,I,v_max,c,T,\ Max\_Size,\ I,\ v\_max,\ c, and Max_ClausesMax\_Clauses may be found in Section 8.1.
6.2 ApproxLearner